Consider a wavelength multiplexer made from


circulators and fiber Bragg gratings. Assume the


input optical power of each wavelength is through


24 is 1 mW. Let both the insertion loss and the


throughput loss of each FBG be 0.25 dB, and let the


insertion loss of a circulator be 0.6 dB. What is the


power level of each wavelength in through h4


emerging from the final circulator ?


1
Expert's answer
2021-01-13T11:44:14-0500

"P_m=(P_o)4+10\log \left(\frac{W}{T}\right)"

"P_o=\log 24"

"P_o=1.380"

"P_m=(1.380)4+10\log \left(\frac{0.25}{0.6}\right)=1.718mW"